# Matrix models as tau functions of the UC hierarchy of integrable equations. II

### v1

@@ Abstract
Keywords:matrix models, partition functions, universal character, tau function, embedded graphs
@@ 1 Introduction
More than half a century passed from the basic discoveries concerning wonderful properties of integrability which were resumed in the famous book [1] edited by Sergei Novikov (see also the book [2] with the Russian translation edited by Sergei Novikov). Since then, the topic was developed in many directions. One of these directions was an application to theory of random matrices [3, 4, 5], which possesses a lot of links with various problems of mathematics, physics, statistics and transmission of information. Here we describe a relation between one of the modification of integrable hierarchies called UC hierarchy [6] and a few new families of solvable matrix models. “Solvable” means that the related integrals can be written as tau function. UC tau functions introduced in [6] can be written in the form of special series of the so-called universal characters (UC) associated with pairs of partitions.
@@
Matrix models can be introduced as the mathematical expectation value of exponentials (or, more generally, the mathematical expectation value of tau functions), the exponents of the exponentials being functions of certain matrices, which are assumed to be random variables, and also functions of a set of parameters. The parameters can be either sets of numbers, which we call coupling constants, or of constant matrices, which we call source matrices. The mathematical expectation is defined within the framework of given matrix ensembles. Thus, the new families of matrix models that we introduce in this paper are tau functions, which are defined either as functions of coupling constants (then the coupling constants play the role of times), or as functions of source matrices (then one obtains tau functions in Miwa variables). We use the expansion of matrix integrals into the above-mentioned universal characters. These characters can respectively be functions of times, or can be functions of source matrices, which are two different cases that our further consideration is based on.
@@
which is independent of $N^{\prime}$ if we put $s_{(k)}=0$ for $k<0$. Hereafter, we denote through $l(\lambda$ the length of the partition $\lambda$.
@@
We use the same notion of $s_{\lambda}$ for the Schur functions as functions of different variables: both $\mathbf{p}$ and $x_{1},\dots,x_{N}$, hopefully this will not cause confusion.
@@
In particular, $s_{\lambda/0}=s_{\lambda}$.
Now, consider two sets of variables $\mathbf{p}$ and $\mathbf{p}^{*}$ where $\deg p_{n}=n$ and $\deg p^{*}_{n}=-n$. The universal character is labeled by a pair of partitions and can be defined as the following quasi-homogeneous polynomial
[…]
@@
The notion of the universal character originally appeared as a generalization of the rational representation of the linear group, see [8] and [6]. In this case, $X$ and $Y$ are related: $Y=X^{-1}\in\mathbb{GL}_{N}$.
@@
where $s_{\mu/\nu}$ are the skew Schur functions [7] and also as
@@
Formula (9) provides a procedure to build the universal character from a pair of Schur functions, we call such a procedure of getting the universal character mating of a pair of the Schur functions $[s_{\mu}\ast s_{\lambda}]=s_{[\mu,\lambda]}$.
@@
We define three ways of mating $[f\ast g]$ of a pair of functions $f$ and $g$, denoted by $[f\ast g]_{1}$, $[f\ast g]_{2}$ and $[f\ast g]_{3}$, where the first two of them are related to the functions of power sum variables $f=f(\mathbf{p}),\,g=g(\mathbf{p}^{*})$, and the last one mates functions of matrices $f=f(X),\,g=g(Y)$:
@@
    In particular $[s_{\mu}(\mathbf{p})\ast s_{\lambda}(\mathbf{p}^{*})]_{1}=[s_{\lambda}(\mathbf{p}^{*})\ast s_{\mu}(\mathbf{p})]_{1}=s_{[\lambda,\mu]}(\mathbf{p},\mathbf{p}^{*})$.
@@
    $[f\ast g]_{2}$:
    $[f(\mathbf{p})\ast g(\mathbf{p}^{*})]_{2}:=\int_{\mathbb{U}_{N}}f\left(\mathbf{p}+\mathbf{p}(U)\right)g\left(\mathbf{p}^{*}-\mathbf{p}(U^{\dagger})\right)d_{*}U\equiv$ (12)
    $[g(\mathbf{p}^{*})\ast f(\mathbf{p})]_{2}=\int_{\mathbb{U}_{N}}f\left(\mathbf{p}-\mathbf{p}(U^{\dagger})\right)g\left(\mathbf{p}^{*}+\mathbf{p}(U)\right)d_{*}U,$ (13)
    where the notation $\mathbf{p}(U)$ was defined in (6). In particular,
@@
    $=\int_{\mathbb{U}_{N}}\sum_{\nu}s_{\mu/\nu}(\mathbf{p})s_{\nu}(U)\sum_{\alpha}s_{\lambda/\alpha}(\mathbf{p}^{*})s_{\alpha}\left(-\mathbf{p}(U^{\dagger})\right)d_{*}U=$ (15)
    $=\sum_{\nu,\alpha}s_{\mu/\nu}(\mathbf{p})s_{\lambda/\alpha}(\mathbf{p}^{*})(-1)^{|\alpha|}\int_{\mathbb{U}_{N}}s_{\nu}(U)s_{\alpha^{t}}(U)d_{*}U$ (16)
    $=\sum_{\nu\atop\ell(\nu),\ell(\nu^{t})\leq N}(-1)^{|\nu|}s_{\mu/\nu}(\mathbf{p})s_{\lambda/\nu^{t}}(\mathbf{p}^{*}),$ (17)
@@
    One can also see that $[s_{\mu}(\mathbf{p})\ast s_{\lambda}(\mathbf{p}^{*})]_{2}=[s_{\lambda}(\mathbf{p})\ast s_{\mu}(\mathbf{p}^{*})]_{2}$.
@@
[…]
@@
In the Miwa variables, it is $\tau^{KP}(X)=\sum_{\lambda}\pi_{\lambda}s_{\lambda}(\mathbf{p}(X))=\sum_{\lambda}\pi_{\lambda}s_{\lambda}(X)$. Here the set of coordinates $\pi_{\lambda}$ defines a solution to the KP equation. These coordinates solve special bilinear relations: the Plücker relations for the Plücker coordinates of a chosen point of the Sato Grassmannian [9, 13, 14],[15, 16, 17]. Each point of the Sato Grassmannian yields a solution to the infinite set of the higher KP equations (the KP hierarchy) and to the KP equation itself111 If one thinks about the KP equation as the (2+1)-dimensional equation of waves in ocean, then, roughly speaking, a point of the infinite Sato Grassmannian can be viewed as an infinite matrix which defines the initial configuration of the surface profile which is a function of two variables..
@@
$\tau^{UC}(\tilde{\mathbf{p}},\mathbf{p})=\tilde{\tau}^{KP}(\tilde{\mathbf{p}}-\partial)\cdot\tau^{KP}(\mathbf{p})=\tilde{\tau}^{KP}(\mathbf{p}-\tilde{\partial})\cdot\tilde{\tau}^{KP}(\tilde{\mathbf{p}})=$ (28)
$=\sum_{\lambda,\mu}\tilde{\pi}_{\lambda}\pi_{\mu}s_{[\lambda,\mu]}(\tilde{\mathbf{p}},\mathbf{p}),$ (29)
@@
The Schur functions are the simplest polynomial KP tau function; choosing $s_{\lambda}(\tilde{\mathbf{p}})$ and $s_{\mu}(\mathbf{p})$ as the KP tau functions, one gets the simplest polynomial UC tau function $s_{[\lambda,\mu]}(\tilde{\mathbf{p}},\mathbf{p})$.
Any KP tau function solves the KP Hirota bilinear equation [17], which is a difference equation involving an arbitrary shift, and, because of the grading of the power sum variables, this equation results in the infinite number of differential equations with respect to the variables $p_{1},p_{2},\dots$ commonly called KP hierarchy.
@@
There are few ways to relate the two topics. They are based on different choices of the mating $[f\ast g]_{i},\,i=0,1,2$, see below.
@@
$=\sum_{\ell(\lambda),\ell(\mu)\leq\min(N_{1},N_{2})}s_{\lambda}(\mathbf{p}_{+}^{(1)})s_{[\lambda,\mu]}(\mathbf{p}_{-}^{(1)},\mathbf{p}_{-}^{(2)})s_{\mu}(\mathbf{p}_{+}^{(2)}).$ (34)
@@
Example 2. For $\tau^{KP}_{1}(\tilde{\mathbf{p}})$, one takes the simplest one, $\tilde{\mathbf{p}}=(\tilde{p},0,0,\dots)$, and, for $\tau^{KP}_{2}(\mathbf{p})$, one takes, say, the well-known one-matrix model [19] where $\mathbf{p}=(p_{1},p_{2},\dots)$ are coupling constants. One obtains
@@
where the measure $dM$ is given by (91) in Appendix A. One can treat this method as an application of the mating.
@@
$=s_{[\lambda,\mu]}(\mathbf{p}^{(1)},\mathbf{p}^{(2)}),$ (38)
@@
see below. The present paper deals mainly with relations between matrix models and the universal characters and the UC tau functions given by (39).
@@ 2.1 General description
$=[Z_{1}(\mathbf{p}^{1},\mathbf{p}(X_{1})|\dots)\ast Z_{1}(\mathbf{p}^{2},\mathbf{p}(X_{2})|\dots)]_{i}=\sum_{\lambda,\mu}(\dots)_{i}s_{\lambda}(\mathbf{p}(X_{1}))s_{[\lambda,\mu]}(\mathbf{p}^{1},\mathbf{p}^{2})s_{\mu}(\mathbf{p}(X_{2})),\quad i=1,2.$
@@
$=\sum_{\lambda,\mu}(\dots)s_{\lambda}(\mathbf{p}^{1})s^{[n]}_{[\lambda,\mu]}(\mathbf{p}(X_{1}),\mathbf{p}(X_{2}))s_{\mu}(\mathbf{p}^{2}),$
@@ 2.2.1 Example 1
We denote by $\mathbf{p}$ the four sets $\mathbf{p}^{(a)}=(p^{(a)}_{1},p^{(a)}_{2},\dots)$, $\tilde{\mathbf{p}}^{(a)}=(\tilde{p}^{(a)}_{1},\tilde{p}^{(a)}_{2},\dots)$, $a=1,2$, and the two matrices $X_{a}$, $a=1,2$ play the role of coupling constants of the model. Here $d_{*}U$ (and $d_{*}V$) is the Haar measure on the unitary group (see (90) in Appendix A).
Integral (40) can be evaluated in an explicit way. First, we recall that the Schur function $s_{\lambda}(X),\,X\in{\rm GL}_{N}$ is the character of the polynomial representation of the linear group labeled by the partition $\lambda$. A pair of partitions labels the rational representation of the linear group, and the related character is called universal character [8] and is denoted by $s_{[\lambda,\mu]}(X_{1},X_{2})$. The both characters can be rewritten in terms of the power sum variables. That is, if
@@
where $s_{\lambda}(I_{N}):=s_{\lambda}(\mathbf{p}(I_{N}))$ is known to be dimension of the irreducible representation of the linear group $\mathbb{GL}_{N}$ [7], which can be written down in an explicit way in terms of the partition $\lambda$ :
$\dim_{\mathbb{GL}_{N}}\lambda=s_{\lambda}(I_{N})=\frac{\prod_{i<j\leq N}(\lambda_{i}-\lambda_{j}-i+j)}{\prod_{i=1}^{N}(\lambda_{i}-i+N)!}\prod_{(i,j)\in\lambda}(N+j-i).$ (47)
The proof of Proposition 1 follows from one of the definitions of the universal character (see formula (*) at the bottom of page 70 in [8]):
$s_{[\lambda,\mu]}(\mathbf{p},\mathbf{p}^{*})=\sum_{\nu}(-1)^{|\nu|}s_{\lambda/\nu}(\mathbf{p})s_{\mu/\nu^{t}}(\mathbf{p}^{*}),$ (48)
(where $s_{\lambda/\nu}$ is the skew Schur function, and $\nu^{t}$ is the partition conjugated to $\nu$) and from the following set of lemmas.
@@
which is the Cauchy-Littlewood identity. We apply it to the integrand of (40).
@@
This model is associated with the graph which is a loop drawn on the sphere: one edge, one vertex and two faces (this can be seen as equator on the globe with a vertex on the equator), see Section 3.
@@
Let both $c_{i}\leq n$ are integers, and $z_{i}\in\mathbb{C}$ (where $i=1,2$). Then the matrix integral (40) is the tau function $\tau^{UC}(\mathbf{p},\mathbf{p}^{*})$ of the UC hierarchy, introduced in [6] with the UC time variables
@@
$=\tau^{UC}(\mathbf{p}(X_{1}),\mathbf{p}^{*}(X_{2})|c_{1},c_{2},z_{1},z_{2},\mathbf{p}^{(1)},\mathbf{p}^{(2)})$

### v3

# Tau functions of the UC hierarchy as partition functions of matrix models

@@ Abstract
Keywords: matrix models, GinUE, partition functions, Schur functions, universal character, tau function, embedded graphs, corner matrices
@@ 1 Introduction
More than half a century passed from the basic discoveries concerning wonderful properties of integrability which were resumed in the famous book [1] edited by Sergei Novikov (see also the book [2] with the Russian translation edited by Sergei Novikov). Since then, the topic was developed in many directions. One of these directions was an application to theory of random matrices [3, 4, 5], which possesses a lot of links with various problems of mathematics, physics, statistics and transmission of information. Here we describe a relation between one of the modification of integrable hierarchies called UC hierarchy [6] and a few new families of exactly solvable matrix models. “Exactly solvable” means that the related integrals can be written as a tau function. UC tau functions introduced in [6] can be written in the form of special series of the so-called universal characters (UC) associated with pairs of partitions.
@@
Matrix models can be represented as the mathematical expectation of an exponential:
$\langle\exp\sum_{m}\frac{p_{m}}{m}\mathrm{tr}\left(X^{m}\right)\rangle_{\Omega},$
where $X$ is a matrix belonging to some ensemble of random matrices $\Omega$  [^1] The detailed description of various ensembles of random matrices can be found in [5], and $p_{m}$ are parameters. More general matrix models are constructed as averages over multimatrix ensembles (e.g., see [7]), and it is not exponentials that are averaged, but more complex expressions (e.g., tau functions can be averaged, see Appendix in [8], and also [9]).
The parameters of a matrix model can be either sets of numbers, which we call coupling constants, or sets of constant matrices, which we call source matrices. A well-known example of such a model, depending on the source matrix, is the Kontsevich integral [10].
Thus, the new families of matrix models that we introduce in this paper are tau functions, which are defined either as functions of coupling constants (then the coupling constants play the role of times), or as functions of source matrices (then one obtains tau functions in Miwa variables). We use the expansion of matrix integrals into the above-mentioned universal characters. These characters can respectively be functions of times, or can be functions of source matrices, which are two different cases that our further consideration is based on.
If the set of power sum variables $\mathbf{p}=(p_{1},p_{2},\dots)$ is parametries by a matrix $\cal V$}
$p_{m}=p_{m}({\cal V}):=\mathrm{tr}\left({\cal V}^{m}\right),\quad m=1,2,\dots$ (1)
then, the eigenvalues of $\cal V$} we will call Miwa variables, such terminology is of use in works devoted to classical integrability. When we say a tau function is written in Miwa variables we mean that $\tau(\mathbf{p})=\tau(\mathbf{p}({\cal V}))=:\tau({\cal V})$, we hope it will not produce a misunderstanding, we use capital letters for matrices and bold letters for the collection of power sum variables.
@@
which is independent of $N^{\prime}$ if we put $s_{(k)}=0$ for $k<0$. Hereafter, we denote through $\ell(\lambda)$ the length of the partition $\lambda$, see Appendix A.
@@
We use the same notation for $s_{\lambda}$ for the Schur functions as functions of different variables: both $\mathbf{p}$ and $x_{1},\dots,x_{N}$, hopefully this will not cause confusion.
@@
In particular, $s_{\lambda/0}=s_{\lambda}$. It is supposed that the Young diagram of the partition $\mu$ is a subdiagram of the Young diagram of partition $\lambda$, details see in [11].
Now, consider two sets of variables $\mathbf{p}$ and $\mathbf{p}^{*}$ where $\deg p_{n}=n$ and $\deg p^{*}_{n}=-n$. The universal character is labeled by a pair of partitions $\lambda=(\lambda_{1},\dots,\lambda_{l})$ and $\mu=(\mu_{1},\dots,\mu_{l^{\prime}})$ and can be defined as the following quasi-homogeneous polynomial
$s_{[\lambda,\mu]}({\bf p},{\bf p}^{*})=\mathrm{det}\left(s_{\mu_{l^{\prime}-i+1}+i-j}({\bf p}),\,1\leq i\leq l^{\prime}\atop s_{\lambda_{i-l^{\prime}}+i-j}({\bf p}^{*}),\,l^{\prime}<i\leq l^{\prime}+l\right)_{1\leq i,j\leq l+l^{\prime}},\quad\deg s_{[\mu,\lambda]}=|\mu|-|\lambda|.$ (6)
@@
The notion of the universal character originally appeared as a generalization of the character of the rational representation of the linear group, see [12] and [6]. In this case, $X$ and $Y$ are related: $Y=X^{-1}\in\mathbb{GL}_{N}$.
@@
where $\nu^{t}$ denotes the partition conjugated to $\nu$ (see [11]) and where $s_{\mu/\nu}$ are the skew Schur functions [11], and also as
@@
Formula (10) provides a procedure to build the universal character from a pair of Schur functions, we call such a procedure of getting the universal character mating of a pair of the Schur functions $s_{\mu}\ast s_{\lambda}]=s_{[\mu,\lambda]}$.
@@
We define three ways of mating $f\ast g]$ of a pair of functions $f$ and $g$, denoted by $f\ast g]_{1}$, $f\ast g]_{2}$ and $f\ast g]_{3}$, where the first two of them are related to the functions of power sum variables $f=f(\mathbf{p}),\,g=g(\mathbf{p}^{*})$, and the last one mates functions of matrices f=f(X),g=g(Y)f=f(X),\,g=g(Y):
@@
    In particular $s_{\mu}(\mathbf{p})\ast s_{\lambda}(\mathbf{p}^{*})]_{1}=[s_{\lambda}(\mathbf{p}^{*})\ast s_{\mu}(\mathbf{p})]_{1}=s_{[\lambda,\mu]}(\mathbf{p},\mathbf{p}^{*})$.
@@
    $f\ast g]_{2}$:
    $f(\mathbf{p})\ast g(\mathbf{p}^{*})]_{2}:=\int_{\mathbb{U}_{N}}f\left(\mathbf{p}+\mathbf{p}(U)\right)g\left(\mathbf{p}^{*}-\mathbf{p}(U^{\dagger})\right)d_{*}U$ (13)
    where $d_{*}U$ denotes the Haar measure on the unitary group $\mathbb{U}_{N}$ and where the notation $\mathbf{p}(U)$ was defined in (7). One can prove that $f(\mathbf{p})\ast g(\mathbf{p}^{*})]_{2}=[g(\mathbf{p}^{*})\ast f(\mathbf{p})]_{2}$. In particular,
@@
    $\int_{\mathbb{U}_{N}}\sum_{\nu}s_{\mu/\nu}(\mathbf{p})s_{\nu}(U)\sum_{\alpha}s_{\lambda/\alpha}(\mathbf{p}^{*})s_{\alpha}\left(-\mathbf{p}(U^{\dagger})\right)d_{*}U=$
    $\sum_{\nu,\alpha}s_{\mu/\nu}(\mathbf{p})s_{\lambda/\alpha}(\mathbf{p}^{*})(-1)^{|\alpha|}\int_{\mathbb{U}_{N}}s_{\nu}(U)s_{\alpha^{t}}(U)d_{*}U$
    $\sum_{\nu\atop\ell(\nu),\ell(\nu^{t})\leq N}(-1)^{|\nu|}s_{\mu/\nu}(\mathbf{p})s_{\lambda/\nu^{t}}(\mathbf{p}^{*}),$ (15)
@@
    One can also see that $s_{\mu}(\mathbf{p})\ast s_{\lambda}(\mathbf{p}^{*})]_{2}=[s_{\lambda}(\mathbf{p})\ast s_{\mu}(\mathbf{p}^{*})]_{2}$.
@@
    $\left[f(A_{1})\ast g(A_{2})\right]_{3}=\int_{{\mathbb{U}}_{n}\times{\mathbb{U}}_{n}}dU_{1}dU_{2}\prod_{a=1,2}\mathrm{det}\left(I_{n}\otimes I_{n}-U_{1}\otimes U_{2}\right)f\left(A_{1}(X_{1},U^{\dagger}_{1})\right)g\left(A_{2}(X_{2},U^{\dagger}_{2})\right),$ (17)
@@
In the Miwa variables, it is $\tau^{KP}(X)=\sum_{\lambda}\pi_{\lambda}s_{\lambda}(\mathbf{p}(X))=\sum_{\lambda}\pi_{\lambda}s_{\lambda}(X)$. Here the set of coefficients $\pi_{\lambda}$ defines a solution to the KP equation. These coefficients solve special bilinear relations: the Plücker relations for the Plücker coordinates of a chosen point of the Sato Grassmannian [13, 17, 18],[19, 20, 21]. Each point of the Sato Grassmannian yields a solution to the infinite set of the higher KP equations (the KP hierarchy) and to the KP equation itself222 If one thinks about the KP equation as the (2+1)-dimensional equation of waves in ocean, then, roughly speaking, a point of the infinite Sato Grassmannian can be viewed as an infinite matrix which defines the initial configuration of the surface profile which is a function of two variables..
@@
[…]
$\sum_{\lambda,\mu}\tilde{\pi}_{\lambda}\pi_{\mu}s_{[\lambda,\mu]}(\tilde{\mathbf{p}},\mathbf{p}),$ (27)
@@
[…]
Any KP tau function solves the KP Hirota bilinear equation [21], which is a difference equation involving an arbitrary shift of KP higher times [21], and, because of the grading of the power sum variables, this equation results in the infinite number of differential equations with respect to the variables $p_{1},p_{2},\dots$ commonly called KP hierarchy.
@@
There are few ways to relate the two topics. They are based on different choices of the mating $f\ast g]_{i},\,i=0,1,2$, see below.
@@
$\sum_{\ell(\lambda),\ell(\mu)\leq\min(N_{1},N_{2})}s_{\lambda}(\mathbf{p}_{+}^{(1)})s_{[\lambda,\mu]}(\mathbf{p}_{-}^{(1)},\mathbf{p}_{-}^{(2)})s_{\mu}(\mathbf{p}_{+}^{(2)}).$
@@
$\tau^{UC}(\mathbf{p},\tilde{\mathbf{p}})=\int\left(\tilde{p}-\mathrm{tr}M\right)e^{\sum_{m>0}\frac{1}{m}p_{m}\mathrm{tr}M^{m}}dM,$ (32)
where the measure dMdM is the standard measure on Hermitian matrices [5]. One can treat this method as an application of the mating.
@@
$s_{[\lambda,\mu]}(\mathbf{p}^{(1)},\mathbf{p}^{(2)}),$ (35)
@@
see below.
This paper mainly considers the relationships between matrix models whose partition functions are expressed through universal characters given by (36) and UC tau functions.
@@ 2.1 General description
$Z_{1}(\mathbf{p}^{1},\mathbf{p}(X_{1})|\dots)\ast Z_{1}(\mathbf{p}^{2},\mathbf{p}(X_{2})|\dots)]_{i}=\sum_{\lambda,\mu}(\dots)_{i}s_{\lambda}(\mathbf{p}(X_{1}))s_{[\lambda,\mu]}(\mathbf{p}^{1},\mathbf{p}^{2})s_{\mu}(\mathbf{p}(X_{2})),\quad i=1,2.$
@@
$\sum_{\lambda,\mu}(\dots)s_{\lambda}(\mathbf{p}^{1})s^{[n]}_{[\lambda,\mu]}(\mathbf{p}(X_{1}),\mathbf{p}(X_{2}))s_{\mu}(\mathbf{p}^{2}),$
@@ 2.2.1 Example 1
[…]
Integral (37) can be evaluated in an explicit way. First, we recall that the Schur function $s_{\lambda}(X),\,X\in{\mathbb{GL}}_{N}$ is the character of the polynomial representation of the linear group labeled by the partition $\lambda$. A pair of partitions labels the rational representation of the linear group, and the related character is called universal character [12] and is denoted by $s_{[\lambda,\mu]}(X_{1},X_{2})$. The both characters can be rewritten in terms of the power sum variables. That is, if
@@
where $s_{\lambda}(I_{N}):=s_{\lambda}(\mathbf{p}(I_{N}))$ is known to be dimension of the irreducible representation $\rho_{\lambda}$ of the linear group $\mathbb{GL}_{N}$ lebelled by $\lambda$ [11], which can be written down in an explicit way in terms of the partition $\lambda$ :
$\dim_{\mathbb{GL}_{N}}\rho_{\lambda}=s_{\lambda}(I_{N})=\frac{\prod_{i<j\leq N}(\lambda_{i}-\lambda_{j}-i+j)}{\prod_{i=1}^{N}(\lambda_{i}-i+N)!}\prod_{(i,j)\in\lambda}(N+j-i).$ (44)
The proof of Proposition 1 follows from one of the definitions of the universal character given by formula (9) and from the following set of lemmas:
@@
which is the Cauchy-Littlewood identity [11]. We apply it to the integrand of (37).
@@
This integral is associated with the graph which is a loop drawn on the sphere: one edge, one vertex and two faces (this can be seen as equator on the globe with a vertex on the equator), see Section 3 and Fig 1b in Appendix D.
@@
Let both $c_{i}\leq n$ be integers, and $z_{i}\in\mathbb{C}$ (where $i=1,2$). Then the matrix integral (37) is the tau function $\tau^{UC}(\mathbf{p},\mathbf{p}^{*})$ of the UC hierarchy, introduced in [6] with the UC time variables
@@
$\tau^{UC}(\mathbf{p}(X_{1}),\mathbf{p}^{*}(X_{2})|c_{1},c_{2},z_{1},z_{2},\mathbf{p}^{(1)},\mathbf{p}^{(2)})$